You have a Gold bar that is 10cm x 4cm x 3cm. And is melted down and 2cm squared bars made instead. How many smaller bars would you be able to make?
Answer by Alan3354(69443)

You have a Gold bar that is 10cm x 4cm x 3cm. And is melted down and 2cm squared bars made instead. How many smaller bars would you be able to make?
---------
10*4*3 = 120 cc
------------
If you mean 2 cm cubes:
2*2*2 = 8 cc
=======================
120/8 = 15 cubes